# GTA: A Benchmark for General Tool Agents

### v1

@@ 3.4 Dataset Statistics
GTA comprises a total of 229 questions, with the basic dataset statistics presented in Table 3. The dataset involves 252 images and 14 distinct tools. It includes 156 objective, 16 subjective, and 57 image-generation queries. The number of tools involved in each question varies from 1 to 4, with most questions using 2 or 3 tools. The steps to resolve the questions range from 2 to 8, with most questions requiring 2 to 4 steps, as depicted in Figure 3(a). The detailed frequency distribution of different tool combinations is listed in Figure 3(b). P, O, L, C are short for Perception, Operation, Logic, Creativity, respectively. Perception+Logic and Perception+Operation are the most frequently appearing tool combination types.
@@ 4.1 Experiment Settings
We evaluate 16 LLMs on GTA. For API-based models, we select GPT-3.5 [19], GPT-4 [1], GPT-4o, Claude3 [2], and Mistral-large [8]. For open-source models, we select LLaMA3 [14] series, Qwen1.5 [3] series, Mistral [8], Mixtral [9], Yi [31] series, Deepseek [4] series. Experiments are conducted using NVIDIA A100 GPU within OpenCompass [6] evaluation platform. We adopt Lagent [25] as the agent framework. ReAct [30] is used as the tool invocation prompt schema. More experiment information can be found in Appendix D.1 and D.2.
We evaluate the models in two modes. Step-by-step mode is designed to evaluate the model’s fine-grained tool-use capabilities. In this mode, the model is provided with the initial $n$ steps of the reference tool chain as prompts, with the expectation to predict the action in step $n+1$. This method does not involve the actual use of the tool, and the prediction of each step does not depend on the model’s preceding outputs. This enables an alignment comparison between the model’s output with each step of the ground truth tool chain. End-to-end mode is designed to reflect the tool agent’s actual task executing performance. In this mode, the model actually calls the tools and solves the problem by itself. Each step relies on the preceding step’s output. We compare the tools selected and the execution result with the ground-truth tool set and the ground-truth result under this mode.
@@ 4.2 Evaluation Metrics
In calculating the metric AnsAcc, we exclude image generation queries and focus solely on queries with pure text answers, including objective and subjective queries. For objective queries, the ground truth contains both a whitelist and a blacklist of phrases. An answer is considered correct if it includes all terms from the whitelist and excludes all terms from the blacklist. In the case of subjective queries, the ground truth contains three manually labeled responses from distinct annotators. We compute the cosine similarity (ranging from 0 to 1) between the model’s prediction and each of the three ground truth answers, ultimately considering the highest score obtained.
Table 4: Main results of GTA. Inst., Tool., Arg., Summ., and Ans. denote InstAcc, ToolAcc, ArgAcc SummAcc, and AnsAcc, respectively. P., O., L., C. denote the F1 score of tool selection in Perception, Operation, Logic, and Creativity categories. Bold denotes the best score among all models. Underline denotes the best score under the same model scale. AnsAcc reflects the overall performance.
@@
Inst. Tool. Arg. Summ. P. O. L. C. Ans.
@@ 4.3 Main Results
Real-world tool-use tasks are challenging for existing LLMs. Current LLMs are struggling to accurately invoke tools to solve these real-world tasks. As shown in Table 4, the best-performing models, GPT-4 and GPT-4o can only correctly solve fewer than 50% of the problems, while the rest of the models solve less than 25%. This shows that real-world problems with implicit steps, real tool invocations, and multimodal contextual inputs impose high requirements on the tool-use capabilities of LLMs. Regarding model performance comparisons, API-based models outperform open-source ones. Among open-source models, Qwen-72b has the highest result accuracy. Larger models within the same series perform better than their smaller counterparts, but larger models from different series do not necessarily outperform the smaller ones, as shown in Figure 4. For example, the AnsAcc of LLaMA3-70b is higher than that of LLaMA3-8b, but lower than Qwen-7b. Some prediction examples of different models can be found in Appendix D.3.
@@
The four metrics in the step-by-step mode follow the buckets effect. From the results, we observe that the overall performance of the system is affected by the lowest metric. To verify this observation, we calculate the Pearson correlation coefficients between four metrics (InstAcc, ToolAcc, ArgAcc, SummAcc) and AnsAcc, the result is shown in Figure 5. We find that the correlation coefficient for ArgAcc with AnsAcc is the highest. ArgAcc is low for most models, indicating that the four metrics follow the buckets effect. For example, the scores of LLaMA3-70b in InstAcc, ToolAcc, and SummAcc are higher than those of Qwen1.5-14b, but its ArgAcc is lower than Qwen1.5-14b, resulting in a lower final answer accuracy. The scores of GPT-4o in InstAcc and ToolAcc are higher than GPT-4, but its weaker argument prediction capability leads to a lower accuracy rate in the final result. The reason for the buckets effect is that under our evaluation framework, the model needs to follow user instructions, invoke tools multiple times in the correct format, and summarize the answer based on the returned results. Any error in this process can lead to an incorrect conclusion. Currently, argument prediction is the weakest capability for most models, suggesting that to enhance their general tool-use capabilities, researchers can focus on argument prediction capabilities. This concerns both the value and the format correctness of an argument.
Different series of LLMs exhibit distinct behavioral patterns. We count the number of successful and failed tool calls, illustrated in Figure 6. Successful means there are not any errors in the tool call. GPT-4o has the highest number of successful tool calls, while GPT4 has the highest successful tool call rate. We find that models from different series exhibit distinct behavioral tendencies. Yi and Deepseek series tend to be aggressive, leaning towards invoking tools frequently but lacks sufficient instruction-following ability to invoke tools in a correct format. The Qwen series is conservative, preferring to invoke tools less often, yet it has stronger instruction-following capabilities than most other open-source models, resulting in a higher success rate of tool calls. The GPT series is neutral, tending to invoke tools moderately and possessing robust instruction-following abilities, which leads to the highest final answer accuracy. This suggests that to improve the performance of Yi or Deepseek, focus should be given to enhancing their instruction-following ability. Conversely, to enhance the
@@
llama3-8b-instruct 20.47 65.15 14.38
llama3-70b-instruct 29.51 69.7 0.8
@@
Qwen series, reducing its conservative behavior to tool invocation could be beneficial.
@@ 5 Conclusion
We propose GTA, a real-world tool-use benchmark for general-purpose agents. The user queries are human-designed, step-implicit, and settled in real-world scenarios. Multimodal contextual inputs are provided. We build an evaluation platform equipped with executable tools in the categories of perception, operation, logic, and creation. Fine-grained metrics are designed for the tool-use capabilities of LLMs in real-world scenarios. We evaluate the tool-use capabilities of 16 LLMs. The evaluation results show that GTA is challenging for current LLMs, with advanced models like GPT-4 struggling with these real-world tasks, completing less than 50% of them. Based on our findings, we give takeaways and further suggestions on tool-use capability improvement for different models. We believe that the GTA benchmark will advance further research in identifying the model’s tool-use capabilities and contribute to realizing general-purpose tool agents.
@@ A.1 Motivation
    This project is supported by the National Key R&D Program of China (No. 2022ZD0161600).
@@ B.2 Examples of Three Query Types
The examples of objective queries $\mathcal{Q}_{o}$, subjective queries $\mathcal{Q}_{s}$, and image generation queries $\mathcal{Q}_{g}$ are shown in Figure 7 to 12, Figure 13 to 16, and Figure 17 to 21, respectively. We provide the complete data sample, which is in the JSON format, including the involved tools, files, query, tool chain, and the final answer. To facilitate automatic evaluation, we design different final answer format for the three query types. For objective queries, the final answer contains both a whitelist and a blacklist of phrases, as shown in Figure 12. An answer is considered correct if it includes all terms from the whitelist and excludes all terms from the blacklist. In the case of subjective queries, the final answer contains three manually labeled responses from distinct annotators, as shown in Figure 16. We compute the cosine similarity (ranging from 0 to 1) between the model’s prediction and each of the three ground truth answers, ultimately considering the highest score obtained. For image generation queries, the final answer is none, as shown in Figure 21, since we do not evaluate the generated images.
@@
Refer to caption Steps: 1. Count the number of eggs in the photo. 2. Identify the eggs needed for one serving of a dish on the recipe. 3. Calculate how many eggs are needed for 12 dishes. 4. Calculate how many boxes of eggs are needed. Answer: 2 Query and Tool Chain in JSON Format: {"tools": [ { "name": "ImageDescription", "description": "A useful tool that returns a brief description of the input image.", "inputs": [ { "type": "image", "name": "image", "description": null, "optional": false, "default": null, "filetype": null } ], "outputs": [ { "type": "text", "name": null, "description": null,
Figure 7: (1 / 6) An example of objective query $\mathcal{Q}_{o}$. The final answer is a uniquely determined number or phrase.
"optional": false, "default": null, "filetype": null } ] }, { "name": "OCR", "description": "This tool can recognize all text on the input image.", "inputs": [ { "type": "image", "name": "image", "description": null, "optional": false, "default": null, "filetype": null } ], "outputs": [ { "type": "text", "name": null, "description": "OCR results, include bbox in x1, y1, x2, y2 format and the recognized text.", "optional": false, "default": null, "filetype": null } ] }, { "name": "CountGivenObject", "description": "The tool can count the number of a certain object in the image.", "inputs": [ { "type": "image", "name": "image", "description": null, "optional": false, "default": null, "filetype": null }, { "type": "text", "name": "text", "description": "The object description in English.",
Figure 8: (2 / 6) An example of objective query $\mathcal{Q}_{o}$. The final answer is a uniquely determined number or phrase.
"optional": false, "default": null, "filetype": null } ], "outputs": [ { "type": "int", "name": null, "description": null, "optional": false, "default": null, "filetype": null } ] } ], "files": [ { "type": "image", "path": "image/image_9.jpg", "url": "https://www.patersonsfruitandveg.co.uk/15-large_default/free-range-eggs-box-of-6.jpg" }, { "type": "image", "path": "image/image_10.jpg", "url": null } ], "dialogs": [ { "role": "user", "content": " I need to prepare twelve servings of this dish. How many boxes of eggs will I need in total?" }, { "role": "assistant", "tool_calls": [ { "type": "function", "function": { "name": "ImageDescription", "arguments": { "image": "image/image_9.jpg" } } } ],
Figure 9: (3 / 6) An example of objective query $\mathcal{Q}_{o}$. The final answer is a uniquely determined number or phrase.
[…]
Figure 10: (4 / 6) An example of objective query $\mathcal{Q}_{o}$. The final answer is a uniquely determined number or phrase.
[…]
Figure 11: (5 / 6) An example of objective query $\mathcal{Q}_{o}$. The final answer is a uniquely determined number or phrase.
[…]
Figure 12: (6 / 6) An example of objective query $\mathcal{Q}_{o}$. The final answer is a uniquely determined number or phrase.
@@
Steps: 1. Recognize the image background and the icon on the sign. 2. Recognize the text in the picture. Answer: You should avoid swimming due to the dangerous current. Query and Tool Chain in JSON Format: {"tools": [ { "name": "OCR", "description": "This tool can recognize all text on the input image.", "inputs": [ { "type": "image", "name": "image", "description": null, "optional": false, "default": null, "filetype": null } ], "outputs": [ { "type": "text", "name": null, "description": "OCR results, include bbox in x1, y1, x2, y2 format and the recognized text.", "optional": false, "default": null, "filetype": null } ] }, { "name": "ImageDescription", "description": "A useful tool that returns a brief description of the input image.", "inputs": [ { "type": "image", "name": "image", "description": null, "optional": false,
Figure 13: (1 / 4) An example of subjective query $\mathcal{Q}_{s}$. The final answer is usually some descriptive text. It is not unique, but the general idea is the same.
"default": null, "filetype": null } ], "outputs": [ { "type": "text", "name": null, "description": null, "optional": false, "default": null, "filetype": null } ] } ], "files": [ { "type": "image", "path": "image/image_27.jpg", "url": "https://images.sbs.com.au/dims4/default/8a59a3f/2147483647/strip/true/crop/3000x1688+0+0/resize/1280x720!/quality/90/?url=http%3A%2F%2Fsbs-au-brightspot.s3.amazonaws.com%2Fe0%2F64%2F10b664a04c59a03b066c4254880e%2Fbeach-safety-sign-cropped.jpg&imwidth=1280" } ], "dialogs": [ { "role": "user", "content": "According to the sign, what should I avoid to do now? Why?" }, { "role": "assistant", "tool_calls": [ { "type": "function", "function": { "name": "ImageDescription", "arguments": { "image": "image/image_27.jpg" } } } ], "thought": "I should first describe the image for detailed information." },
Figure 14: (2 / 4) An example of subjective query $\mathcal{Q}_{s}$. The final answer is usually some descriptive text. It is not unique, but the general idea is the same.
{ "role": "tool", "name": "ImageDescription", "content": { "type": "text", "content": "The image features a person standing on a beach, with a yellow sign in the foreground. The sign is placed on the sand, and it appears to be a warning or a caution sign. The person is walking past the sign, possibly enjoying the beach atmosphere. The scene captures a moment of leisure and outdoor activity." } }, { "role": "assistant", "tool_calls": [ { "type": "function", "function": { "name": "OCR", "arguments": { "image": "image/image_27.jpg" } } } ], "thought": "There seems a sign in the image. I will extract all text on the sign." }, { "role": "tool", "name": "OCR", "content": { "type": "text", "content": "(428, 118, 929, 603) X DANGEROUS CURRENT" } }, { "role": "assistant", "content": "You should avoid going into the water now due to the dangerous current." } ],
Figure 15: (3 / 4) An example of subjective query $\mathcal{Q}_{s}$. The final answer is usually some descriptive text. It is not unique, but the general idea is the same.
"gt_answer": [ "You should avoid swimming. Because there is dangerous current.", "You should avoid swimming in the sea because the signs in the picture indicate that it is a dangerous area and swimming can be dangerous.", "According to the sign, I should avoid to go swimming in the sea. The background of the picture is a sea area, and there is a yellow warning sign with DANGEROUS CURRENT written on it. Additionally, a red cross is marked over the act of swimming, indicating that swimming here is dangerous and prohibited. Therefore, I should avoid swimming in the sea." ] }
Figure 16: (4 / 4) An example of subjective query $\mathcal{Q}_{s}$. The final answer is usually some descriptive text. It is not unique, but the general idea is the same.
@@
Figure 17: (1 / 5) An example of image generation query $\mathcal{Q}_{i}$. The final answer is none since we do not evaluate the generated image.
Query and Tool Chain in JSON Format: {"tools": [ { "name": "OCR", "description": "This tool can recognize all text on the input image.", "inputs": [ { "type": "image", "name": "image", "description": null, "optional": false, "default": null, "filetype": null } ], "outputs": [ { "type": "text", "name": null, "description": "OCR results, include bbox in x1, y1, x2, y2 format and the recognized text.", "optional": false, "default": null, "filetype": null } ] }, { "name": "DrawBox", "description": "A tool to draw a box on a certain region of the input image.", "inputs": [ { "type": "image", "name": "image", "description": null, "optional": false, "default": null, "filetype": null }, { "type": "text", "name": "bbox", "description": "The bbox coordinate in the format of ‘(x1, y1, x2, y2)‘", "optional": false, "default": null, "filetype": null },
Figure 18: (2 / 5) An example of image generation query $\mathcal{Q}_{i}$. The final answer is none since we do not evaluate the generated image.
{ "type": "text", "name": "annotation", "description": "The extra annotation text of the bbox", "optional": true, "default": null, "filetype": null } ], "outputs": [ { "type": "image", "name": null, "description": null, "optional": false, "default": null, "filetype": null } ] } ], "files": [ { "type": "image", "path": "image/image_129.jpg", "url": null } ], "dialogs": [ { "role": "user", "content": "I want to go to the highest-rated restaurant.Please circle it in the map." }, { "role": "assistant", "tool_calls": [ { "type": "function", "function": { "name": "OCR", "arguments": { "image": "image/image_129.jpg" } } } ],
Figure 19: (3 / 5) An example of image generation query $\mathcal{Q}_{i}$. The final answer is none since we do not evaluate the generated image.
[…]
Figure 20: (4 / 5) An example of image generation query $\mathcal{Q}_{i}$. The final answer is none since we do not evaluate the generated image.
"function": { "name": "DrawBox", "arguments": { "bbox": "(548, 582, 851, 628)" } } } ], "content": "Based on the OCR results, the highest-rated restaurant identified is \"Trattoria Emilia\" with a rating of 4.6." }, { "role": "tool", "name": "DrawBox", "content": { "type": "image", "content": "image/dummy_generated_image.jpg" } }, { "role": "assistant", "content": "The image is saved." } ], "gt_answer": null }
Figure 21: (5 / 5) An example of image generation query $\mathcal{Q}_{i}$. The final answer is none since we do not evaluate the generated image.

### v2

@@ 3.4 Dataset Statistics
GTA comprises a total of 229 questions, with the basic dataset statistics presented in Table 3. The dataset involves 252 images and 14 distinct tools. It includes 156 objective, 16 subjective, and 57 image generation queries. The number of tools involved in each question varies from 1 to 4, with most questions using 2 or 3 tools. The steps to resolve the questions range from 2 to 8, with most questions requiring 2 to 4 steps, as depicted in Figure 3(a). The detailed frequency distribution of different tool combinations is listed in Figure 3(b). P, O, L, C are short for Perception, Operation, Logic, Creativity, respectively. Perception+Logic and Perception+Operation are the most frequently appearing tool combination types.
@@ 4.1 Experiment Settings
We evaluate 16 LLMs on GTA. For API-based models, we select GPT-3.5 [20], GPT-4 [1], GPT-4o, Claude-3 [2], and Mistral-Large [9]. For open-source models, we select Llama-3 [15] series, Qwen1.5 [3] series, Mistral [9], Mixtral [10], Yi [34] series, Deepseek [4] series. Experiments are conducted using NVIDIA A100 GPU within OpenCompass [7] evaluation platform. We adopt Lagent [27] as the agent framework. ReAct [33] is used as the tool invocation prompt schema. More experiment information can be found in Appendix D.1 and D.2.
We evaluate the models in two modes. Step-by-step mode is designed to evaluate the model’s fine-grained tool-use capabilities. In this mode, the model is provided with the initial $n$ steps of the reference tool chain as prompts, with the expectation to predict the action in step $n+1$. This method does not involve the actual use of the tool, and the prediction of each step does not depend on the model’s preceding outputs. This enables an alignment comparison between the model’s output with each step of the ground truth tool chain. End-to-end mode is designed to reflect the tool agent’s actual task executing performance dynamically. In this mode, the model actually calls the tools and solves the problem by itself. Each step relies on the preceding step’s output. We compare the tools selected and the execution result with the ground-truth tool set and result under this mode.
@@ 4.2 Evaluation Metrics
In calculating the metric AnsAcc, we exclude image generation queries and focus solely on queries with pure text answers, including subjective and objective queries. For objective queries, the ground truth contains both a whitelist and a blacklist of phrases. An answer is considered correct if it includes all terms from the whitelist and excludes all terms from the blacklist. In the case of subjective queries, the ground truth contains three manually labeled responses from distinct annotators. We compute the cosine similarity (ranging from 0 to 1) between the model’s prediction and each of the three ground truth answers, ultimately considering the highest score obtained. We also design a metric AnsAcc w/ ImgGen, to take image generation queries into account indirectly. Given that the outcome of the image generation is determined solely by the input parameters, we evaluate the accuracy of these parameter predictions. If the predicted parameters are correct, the images produced should align with the specified task objectives. The specific score calculation formulas of subjective and image generation queries are shown in Appendix D.3.
Table 4: Main results of GTA. Inst., Tool., Arg., Summ., Ans., Ans.+I denote InstAcc, ToolAcc, ArgAcc SummAcc, AnsAcc, and AnsAcc w/ ImgGen respectively. P., O., L., C. denote the F1 score of tool selection in Perception, Operation, Logic, and Creativity categories. Bold denotes the best score among all models. Underline denotes the best score under the same model scale. AnsAcc reflects the overall performance.
@@
Inst. Tool. Arg. Summ. P. O. L. C. Ans. Ans.+I
@@ 4.3 Main Results
Real-world tool-use tasks are challenging for existing LLMs. Current LLMs are struggling to accurately invoke tools to solve these real-world tasks. As shown in Table 4, the best-performing models, GPT-4 and GPT-4o can only correctly solve fewer than 50% of the problems, while the rest of the models solve less than 25%. This shows that real-world problems with implicit steps, real tool invocations, and multimodal contextual inputs impose high requirements on the tool-use capabilities of LLMs. Regarding model performance comparisons, API-based models outperform open-source ones. Among open-source models, Qwen1.5-72B-Chat has the highest result accuracy. Larger models within the same series perform better than their smaller counterparts, but larger models from different series do not necessarily outperform the smaller ones, as shown in Figure 4. For example, the AnsAcc of Llama-3-70B-Instruct is higher than that of Llama-3-8B-Instruct, but lower than Qwen1.5-7B-Chat.
@@
The current bottleneck mainly lies on argument prediction. From the results, we observe that the overall performance of the system is affected by the lowest metric. We argue that the four metrics in the step-by-step mode follow the buckets effect. To verify this observation, we calculate the Pearson correlation coefficients between four metrics (InstAcc, ToolAcc, ArgAcc, SummAcc) and AnsAcc, the result is shown in Figure 5. We find that the correlation coefficient for ArgAcc with AnsAcc is the highest. ArgAcc is low for most models, indicating that the four metrics follow the buckets effect. For example, the scores of Llama-3-70B-Instruct in InstAcc, ToolAcc, and SummAcc are higher than those of Qwen1.5-14B-Chat, but its ArgAcc is lower than Qwen1.5-14B-Chat, resulting in a lower final answer accuracy. The scores of GPT-4o in InstAcc and ToolAcc are higher than GPT-4, but its weaker argument prediction capability leads to a lower accuracy rate in the final result. The reason for the buckets effect is that under our evaluation framework, the model needs to follow user instructions, invoke tools multiple times in the correct format, and summarize the answer based on the returned results. Any error in this process can lead to an incorrect conclusion. Currently, argument prediction is the weakest capability for most models, suggesting that to enhance their general tool-use capabilities, researchers can focus on argument prediction capabilities. This concerns both the value and the format correctness of an argument.
Different series of LLMs exhibit distinct behavioral patterns. We count the number of successful and failed tool calls, illustrated in Figure 6. Successful means there are not any errors in the tool call. GPT-4o has the highest number of successful tool calls, while GPT-4 has the highest successful tool call rate. We find that models from different series exhibit distinct behavioral tendencies. Yi and Deepseek series tend to be aggressive, invoking tools frequently but lacks sufficient instruction-following ability to invoke tools in a correct format. The Qwen series is conservative, preferring to invoke tools less often, yet it has stronger instruction-following capabilities, resulting in a higher success rate of tool calls. The GPT series is neutral, tending to invoke tools moderately and possessing robust instruction-following abilities, which leads to the highest final answer accuracy. This suggests that to improve the performance of Yi or Deepseek, focus should be given to enhancing their instruction-following ability.
@@
Llama-3-8B-Instruct 20.47 65.15 14.38
Llama-3-70B-Instruct 29.51 69.7 0.8
@@
Conversely, to enhance the Qwen series, reducing its conservative behavior to tool calls could be beneficial.
@@
4.4 Further Analysis and Exploration
Detailed Error Analysis. In Section 4.3, we discuss the bottleneck in task performance arising from most models’ inability to generate responses or predict arguments in the correct format. To understand the reason behind these model failures, we conduct a detailed analysis of the predictions generated by GPT-4-1106-Preview and Llama-3-8B-Instruct. We systematically categorize seven primary error types. The statistical outcomes are presented in Table 6. Detailed error cases of each type can be found in Appendix D.4.
Table 6: Detailed error distribution of GPT-4-1106-Preview and Llama-3-8B-Instruct on argument prediction.
Error Type GPT-4-1106-Preview Llama-3-8B-Instruct
Planning error:
No action, requesting more information from the user. 24 (38.7%) 0
No action, the whole response is model thought. 31 (50%) 2 (0.1%)
Format error:
The arguments do not follow the correct JSON format. 5 (8.1%) 118 (45.4%)
Trying to call multiple tools in one step. 1 (1.6%) 43 (16.5%)
Generating redundant information that leads to incorrect argument parsing. 0 51 (19.6%)
Repeating contents from the prompt. 0 41 (15.8%)
Generating the final answer but not following the correct format. 1 (1.6%) 5 (1.9%)
Total number 62 (100%) 260 (100%)
Our analysis reveals distinct error distributions between GPT-4 and Llama-3. GPT-4 consistently adheres to the given prompts when executing actions, in contrast to Llama-3, which often fails to maintain the prescribed format. However, GPT-4 is prone to generating passive thought processes or attempting interaction with the user, rather than taking decisive action.
For the GPT-4 model, the predominant type of error is No Action, wherein the model neither utilizes tools nor produces a final answer. In 38.7% of erroneous responses, GPT-4 attempts to engage with the user, mistakenly assuming the query lacks clarity and requesting additional information, despite the query and input images supplying sufficient details for task resolution. Furthermore, 50% of the error responses consist solely of the model’s internal thought without any corresponding action.
For the Llama-3 model, most errors are related to formatting during action sequences, such as invoking tools or generating the final answer. Specifically, 45.4% of errors originate from argument predictions not adhering to a valid JSON format. Additionally, in 16.5% of the flawed responses, the model attempts to invoke multiple tools simultaneously, which is not supported by the agent system. Moreover, 19.6% of the errors occur when the model disregards the prompt and generates redundant information after argument prediction, leading to incorrect argument parsing. Finally, in 15.8% of the cases, the model fails to perform the correct action, merely repeating content from the prompt.
Further Exploration to Enhance Model Performance. Since the LLM functions as the central controller of an agent system, producing responses that strictly comply with the agent protocol is important. Fine-tuning on ReAct and JSON format may mitigate format-related errors during action execution. To verify this, we further compare Llama-2-Chat-7B with Agent-Flan-7B on GTA benchmark. AgentFLAN[6] is a popular instruction tuning method that fine-tunes LLM-based agents using ReAct and JSON instruction-following data. Agent-Flan-7B is fine-tuned from Llama-2-Chat-7B using AgentFLAN method. The results are shown in Table 7.
We discovered that Agent-Flan-7B’s InstAcc and ToolAcc metrics were significantly higher than those of Llama-2-Chat-7B. The responses of Agent-Flan-7B follow the format of ’Thought-Action-Action Input’ that specified in the prompt. But most responses of Llama-2-Chat-7B fail to follow the format. This further suggests the improved instruction following capability of Agent-Flan-7B.
However, we note that the ArgAcc of Agent-Flan-7B is still low. We compare the response of the two models, as shown in Appendix D.5. We find that although Agent-Flan-7B follows the format of ’Thought-Action-Action Input’, it sometimes fails to generate the argument (Action Input) in a correct JSON format, or summarizes the final answer incorrectly. Thus, how to further enhance the model’s capabilities on GTA through instruction fine-tuning is still an open problem.
Table 7: Comparison of Llama-2-Chat-7B with Agent-Flan-7B (which is fine-tuned from Llama-2-Chat-7B on ReAct and JSON format data) on GTA.
Model Inst. Tool. Arg. Summ. Ans. Ans.+I
Llama-2-Chat-7B 30.86 16.34 0.36 47.69 3.96 2.98
Agent-Flan-7B 71.60 41.11 6.82 52.88 6.45 4.85
@@ 5 Conclusion
We propose GTA, a real-world tool-use benchmark for general-purpose agents. The user queries are human-designed, step-implicit, and settled in real-world scenarios. Multimodal contextual inputs are provided. We build an evaluation platform equipped with executable tools in the categories of perception, operation, logic, and creation. Fine-grained metrics are designed for the tool-use capabilities of LLMs in real-world scenarios. We evaluate the tool-use capabilities of 16 LLMs. The evaluation results show that GTA is challenging for current LLMs, with advanced models like GPT-4 struggling with these real-world tasks, completing less than 50% of them. Based on our findings, we give takeaways and further suggestions on tool-use capability improvement. We believe that the GTA benchmark will advance further research in identifying the model’s tool-use capabilities and contribute to realizing general-purpose tool agents.
@@ 6 Limitations
7 Acknowledgements
This work is supported by the National Key R&D Program of China (No. 2022ZD0161600), and the National Natural Science Foundation of China under Grants 62422311 and 62176152.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Introduction.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Limitation.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] Our paper proposes a dataset to evaluate LLM-based tool agents in real-world scenarios. There is currently no negative social impact.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] See supplementary materials.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] https://github.com/open-compass/GTA
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] https://github.com/open-compass/GTA
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 4.1.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [N/A]
    2. (b)
      Did you mention the license of the assets? [N/A]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [N/A]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See supplementary materials.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ A.1 Motivation
    This work is supported by the National Key R&D Program of China (No. 2022ZD0161600), and the National Natural Science Foundation of China under Grants 62422311 and 62176152.
@@ A.5 Uses
  • Are there any potential negative social impacts?
    The GTA benchmark may have potential negative societal impacts. These include copyright concerns related to image data collection. The presence of images involving people in our dataset also raises privacy concerns. Additionally, during the evaluation of GTA, the agent system could potentially experience hallucinations and generate harmful information. Besides, given the inclusion of coding questions in GTA, the agent system might produce malicious code.
@@ B.2 Examples of Three Query Types
The examples of objective queries $\mathcal{Q}_{o}$, subjective queries $\mathcal{Q}_{s}$, and image generation queries $\mathcal{Q}_{g}$ are shown in Figure 7, Figure 8, and Figure 9, respectively. In the supplementary material, we provide the complete data sample, which is in the JSON format, including the involved tools, files, query, tool chain, and the final answer. To facilitate automatic evaluation, we design different final answer format for the three query types. For objective queries, the final answer contains both a whitelist and a blacklist of phrases. An answer is considered correct if it includes all terms from the whitelist and excludes all terms from the blacklist. In the case of subjective queries, the final answer contains three manually labeled responses from distinct annotators. We compute the cosine similarity (ranging from 0 to 1) between the model’s prediction and each of the three ground truth answers, ultimately considering the highest score obtained. For image generation queries, the final answer is none, since we evaluate the execution accuracy through measuring the argument accuracy of image generation tools.
@@
Figure 7: An example of objective query $\mathcal{Q}_{o}$. The final answer is a uniquely determined number or phrase.
@@
Steps: 1. Recognize the image background and the icon on the sign. 2. Recognize the text in the picture. Answer: You should avoid swimming due to the dangerous current.
Figure 8: An example of subjective query $\mathcal{Q}_{s}$. The final answer is usually some descriptive text. It is not unique, but the general idea is the same.
@@
Figure 9: An example of image generation query $\mathcal{Q}_{i}$. The final answer is none since we do not evaluate the generated image directly.